# A game-theoretic probability approach to loopholes in CHSH experiments

### v2

@@ 1 Introduction
While most analyses of the CHSH inequality rely on probabilistic or measure-theoretic formulations, comparatively little attention has been paid to how the availability and timing of information themselves constrain possible correlations. The objective of this paper is to shed new light on the conditions under which the CHSH inequality holds or is violated, from this informational perspective. To this end, this paper employs the framework of game-theoretic probability theory [13].
@@ 4.1 A hidden-variable game with the locality loophole
  Hidden-variable game with the locality loophole   Players : Scientist A, Scientist B, Nature A, and Nature B. Protocol : $K_{0}^{A}=K_{0}^{B}=1$. FOR $n\in\mathbb{Z}_{>0}$ : Scientist A announces $\{q_{n}^{A}(\,\cdot\,\mid s)\in\mathcal{P}(\Omega)\}_{s\in\Theta^{A}}$ and $s_{n}\in\Theta^{A}$. Nature A announces $\omega_{n}^{s_{n}}\in\Omega$. Scientist B announces $\{q_{n}^{B}(\,\cdot\,\mid\omega_{n}^{s_{n}},s_{n},t)\in\mathcal{P}(\Omega)\}_{t\in\Theta^{B}}$ and $t_{n}\in\Theta^{B}$. Nature B chooses $\lambda_{n}\in X_{s_{n}}^{-1}(\{\omega_{n}^{s_{n}}\})$ and announces $\omega_{n}^{t_{n}}:=X_{t_{n}}(\lambda_{n})$. $K_{n}^{A}:=K_{n-1}^{A}\cdot\frac{q_{n}^{A}(\omega_{n}^{s_{n}}\mid s_{n})}{p(\omega_{n}^{s_{n}}\mid s_{n})}$, $K_{n}^{B}:=K_{n-1}^{B}\cdot\frac{q_{n}^{B}(\omega_{n}^{t_{n}}\mid\omega_{n}^{s_{n}},s_{n},t_{n})}{p(\omega_{n}^{t_{n}}\mid\omega_{n}^{s_{n}},s_{n},t_{n})}$. END FOR.
@@
In the hidden-variable game with the locality loophole, Scientists A and B have a joint strategy that ensures $\lim_{n\to\infty}K_{n}^{A}=\infty$ or $\lim_{n\to\infty}K_{n}^{B}=\infty$ unless, for any $s\in\Theta^{A}$, $t\in\Theta^{B}$ and $a,b\in\Omega$,
@@ 4.2 A hidden-variable game with loopholes closed
$T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4};s,t):=$ $\left|\left\{k\middle|\begin{aligned} \;&1\leq k\leq n,\,s_{k}=s,\,t_{k}=t,\\ &(X_{1}(\lambda_{k}),X_{2}(\lambda_{k}),X_{3}(\lambda_{k}),X_{4}(\lambda_{k}))=(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\end{aligned}\right\}\right|,$
$T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4}):=$ $\left|\left\{k\middle|1\leq k\leq n,\,(X_{1}(\lambda_{k}),X_{2}(\lambda_{k}),X_{3}(\lambda_{k}),X_{4}(\lambda_{k}))=(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\right\}\right|$
$=$ $\sum_{(s,t)\in\Theta}T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4};s,t),$
$T_{n}(s,t):=$ $|\{k\mid 1\leq k\leq n,\,s_{k}=s,\,t_{k}=t\}|$
$=$ $\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4};s,t).$
The quantity $T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4};s,t)$ counts how often the hidden variable assignments produce the outcome quadruple $(\omega^{1},\omega^{2},\omega^{3},\omega^{4})$ under the measurement setting $(s,t)$. The quantities $T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})$ and $T_{n}(s,t)$ are the corresponding marginal counts. We then define empirical distributions on $\Omega^{4}\times\Theta$, $\Omega^{4}$, and $\Theta$ by
$\hat{P}_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4};s,t)$ $:=\frac{T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4};s,t)}{n},$
$\hat{Q}_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})$ $:=\frac{T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n},$
$\hat{R}_{n}(s,t)$ $:=\frac{T_{n}(s,t)}{n}.$
@@
  Hidden-variable game with loopholes closed   Players : Scientist AB, Nature AB. Protocol : $K_{0}^{AB}=1$, $W_{0}^{AB}=1$. FOR $n\in\mathbb{Z}_{>0}$ : Scientist AB announces $\{(q_{n}^{AB}(\,\cdot,\cdot\,\mid s,t))\in\mathcal{P}(\Omega^{2})\}_{(s,t)\in\Theta}$ and $(s_{n},t_{n})\in\Theta$. Nature AB announces $\lambda_{n}\in\Lambda$ and $(\omega_{n}^{s_{n}},\omega_{n}^{t_{n}}):=(X_{s_{n}}(\lambda_{n}),X_{t_{n}}(\lambda_{n}))$. $K_{n}^{AB}:=K_{n-1}^{AB}\cdot\frac{q_{n}^{AB}(\omega_{n}^{s_{n}},\omega_{n}^{t_{n}}\mid s_{n},t_{n})}{p(\omega_{n}^{s_{n}},\omega_{n}^{t_{n}}\mid s_{n},t_{n})}$. $W_{n}^{AB}:=W_{n-1}^{AB}\cdot\frac{\hat{P}_{n}(\omega_{n}^{1},\omega_{n}^{2},\omega_{n}^{3},\omega_{n}^{4};s_{n},t_{n})}{\hat{Q}_{n}(\omega_{n}^{1},\omega_{n}^{2},\omega_{n}^{3},\omega_{n}^{4})\cdot\hat{R}_{n}(s_{n},t_{n})}$. END FOR.
The second capital process $W_{n}^{AB}$ is designed to test, in a sequential manner, whether the empirical joint distribution on $\Omega^{4}\times\Theta$ factorizes asymptotically into the product of its marginals. It is defined recursively via a likelihood-ratio-type update between $\hat{P}_{n}$ and $\hat{Q}_{n}\cdot\hat{R}_{n}$, so that any systematic dependence between Nature’s hidden-variable declarations and the chosen measurement settings leads to the divergence of $W_{n}^{AB}$.
Structurally, the present game differs from the locality-loophole game in only two respects. First, Scientist AB introduces the additional capital process $W_{n}^{AB}$ in order to operationally enforce a measurement-independence-like constraint. Second, by changing the timing of the announcements (relative to the locality-loophole game), we reformulate the dependence between the hidden variable $\lambda_{n}$ and the measurement settings $(s_{n},t_{n})$ from a purely temporal dependence into a dependence that is empirically testable through frequencies. Although the timing of moves differs from that in the locality-loophole game, this does not restrict Nature’s ability to generate outcome pairs $(\omega_{n}^{s_{n}},\omega_{n}^{t_{n}})$ as functions of $(s_{n},t_{n})$. The essential difference is that the loophole-closed game makes such dependence detectable within the game, which ultimately leads to the following incompatibility result.
@@
In the hidden-variable game with loopholes closed, Scientist AB has a strategy $q:\Omega^{\ast}\rightarrow\mathcal{P}(\Omega^{2})$, that ensures $\lim_{n\to\infty}K_{n}^{AB}=\infty$ unless, for any $(s,t)\in\Theta$ and $a,b\in\Omega$,
@@
In addition, Scientist AB’s second capital process $W_{n}^{AB}$ diverges, $\lim_{n\to\infty}W_{n}^{AB}=\infty$, unless, for any $(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}$ and $(s,t)\in\Theta$,
@@
Furthermore, Nature AB cannot satisfy (1) and (2) simultaneously. Thus in this game, either $\lim_{n\to\infty}K_{n}^{AB}=\infty$ or $\lim_{n\to\infty}W_{n}^{AB}=\infty$ must occur.
@@
Theorem 4.2 states, in game-theoretic terms, that Scientist AB possesses a winning strategy in the hidden-variable game with loopholes closed. More precisely, Scientist AB can choose a strategy $q:\Omega^{\ast}\to\mathcal{P}(\Omega)$ such that Nature AB cannot keep both capital processes $K_{n}^{AB}$ and $W_{n}^{AB}$ bounded: Nature must either reproduce the CHSH table at the level of empirical conditional frequencies (1) or exhibit setting-independence-like behavior, in the frequency sense (2), but it cannot achieve both simultaneously. Thus, as $n\to\infty$, at least one of $K_{n}^{AB}$ or $W_{n}^{AB}$ necessarily diverges, meaning that Nature inevitably loses in the sense of game-theoretic probability.
At first sight this may seem at odds with the empirical situation: in actual CHSH experiments, Bell-type inequalities are violated in accordance with quantum-mechanical predictions. The point, however, is that the theorem is not merely saying that “a violation occurs.” Rather, it clarifies what must give way within a local hidden-variable (LHV) picture once the relevant loopholes are controlled.
The message of Theorem 4.2 is that one cannot simultaneously reproduce the CHSH table (including quantum correlations) and maintain measurement independence (free choice of settings) within an LHV model. While this is often discussed within a probabilistic framework, our theorem recasts it as a sequential, game-theoretic validation that does not assume an underlying probability space.
Scientists run two capital processes in parallel.
  • $K_{n}^{AB}$ tests whether the empirical conditional frequencies converge to the CHSH table.
  • $W_{n}^{AB}$ tests whether the empirical distribution on $\Omega^{4}\times\Theta$ factorizes asymptotically into the product of its marginal distributions on $\Omega^{4}$ and on $\Theta$, providing a frequency-based analogue of measurement independence.
Note that (2) is weaker than the probabilistic independence assumptions usually imposed in the context of the CHSH inequality. Theorem 4.2 asserts that Nature (the LHV side) cannot satisfy these two convergence requirements simultaneously. Consequently, as $n\to\infty$, at least one of $K_{n}^{AB}$ or $W_{n}^{AB}$ must diverge.
Physically, an LHV model can survive asymptotically only by abandoning one of the two: either it fails to reproduce the CHSH table, or it gives up measurement independence by correlating the setting choices with the hidden variable. Operationally, the two capital processes function as sequential test statistics that diagnose compatibility with each requirement along the data stream. Thus, in a loophole-closed setting, Scientists obtain an operational winning strategy: they inevitably detect a breakdown, either in reproducing the CHSH table at the level of empirical frequencies or in setting-independence-like behaviour, thereby excluding LHV explanations in a data-driven sense.
@@ A Proofs of Theorems presented in Section 4
We divide the proof into three parts. First, as a direct consequence of Theorem 2.1, Scientist AB has a strategy $q:\Omega^{\ast}\rightarrow\mathcal{P}(\Omega)$, that ensures $\lim_{n\to\infty}K_{n}^{AB}=\infty$ unless, for all $(s,t)\in\Theta$ and $a,b\in\Omega$,
@@
We note that as in Theorem 4.1, this step is merely an application of Theorem 2.1, which forces Nature to conform to the CHSH table at the level of empirical frequencies. The novelty of Theorem 4.2 lies not in this enforcement itself, but in the fact that its consequence is incompatible with the boundedness of the capital process $W_{n}^{AB}$.
Second, we evaluate the capital process $W_{n}^{AB}$. To simplify the symbols, we write $\tau=(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}$ and $u=(s,t)\in\Theta$. Then by the recursion formula of $W_{n}^{AB}$,
$W_{n}^{AB}=\frac{n!\cdot\prod_{\tau\in\Omega^{4}}\prod_{u\in\Theta}T_{n}(\tau;u)!}{\prod_{\tau\in\Omega^{4}}T_{n}(\tau)!\cdot\prod_{u\in\Theta}T_{n}(u)!}.$
@@
$\frac{\log W_{n}^{AB}}{n}=$ $\frac{1}{n}\left\{\log n!+\sum_{\tau\in\Omega^{4}}\sum_{u\in\Theta}\log T_{n}(\tau;u)!-\sum_{\tau\in\Omega^{4}}\log T_{n}(\tau)!-\sum_{u\in\Theta}\log T_{n}(u)!\right\}$
$=$ $\log n+\sum_{\tau,u}\frac{T_{n}(\tau;u)}{n}\log T_{n}(\tau;u)$
$-\sum_{\tau}\frac{T_{n}(\tau)}{n}\log T_{n}(\tau)-\sum_{u}\frac{T_{n}(u)}{n}\log T_{n}(u)+O\left(\frac{\log n}{n}\right)$
$=$ $\sum_{\tau,u}\frac{T_{n}(\tau;u)}{n}\left(\log\frac{T_{n}(\tau;u)}{n}-\log\frac{T_{n}(\tau)}{n}-\log\frac{T_{n}(u)}{n}\right)+O\left(\frac{\log n}{n}\right)$
@@
then $\limsup W_{n}^{AB}=\infty$.
@@
$\frac{T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4};s,t)}{T_{n}(s,t)}-\frac{T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}$
@@
$\frac{S_{n}(\omega^{s},\omega^{t})}{n}-\frac{S_{n}(\omega^{s},\omega^{t};s,t)}{T_{n}(s,t)}=o(1),$ (A.2)
@@
$S_{n}(\omega^{s},\omega^{t})$ $:=|\{k\mid 1\leq k\leq n,\,X_{s}(\lambda_{k})=\omega^{s},\,X_{t}(\lambda_{k})=\omega^{t}\}|,$
$S_{n}(\omega^{s},\omega^{t};s,t)$ $:=|\{k\mid 1\leq k\leq n,\,X_{s}(\lambda_{k})=\omega^{s},\,X_{t}(\lambda_{k})=\omega^{t},\,s_{k}=s,\,t_{k}=t\}|.$
@@
$C_{n}(s,t):=\sum_{(\omega^{s},\omega^{t})\in\Omega^{2}}\omega^{s}\omega^{t}\cdot\frac{S_{n}(\omega^{s},\omega^{t};s,t)}{T_{n}(s,t)}.$
@@
$=\sum_{(\omega^{1},\omega^{3})\in\Omega^{2}}\omega^{1}\omega^{3}\cdot\frac{S_{n}(\omega^{1},\omega^{3};1,3)}{T_{n}(1,3)}-\sum_{(\omega^{1},\omega^{4})\in\Omega^{2}}\omega^{1}\omega^{4}\cdot\frac{S_{n}(\omega^{1},\omega^{4};1,4)}{T_{n}(1,4)}$
$=\sum_{(\omega^{1},\omega^{3})\in\Omega^{2}}\omega^{1}\omega^{3}\cdot\frac{S_{n}(\omega^{1},\omega^{3})}{n}-\sum_{(\omega^{1},\omega^{4})\in\Omega^{2}}\omega^{1}\omega^{4}\cdot\frac{S_{n}(\omega^{1},\omega^{4})}{n}+o(1)$
$=\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}\omega^{1}\omega^{3}\cdot\frac{T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}-\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}\omega^{1}\omega^{4}\cdot\frac{T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}+o(1)$
$=\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}\omega^{1}\omega^{3}(1\pm\omega^{2}\omega^{4})\cdot\frac{T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}$
$\ \ -\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}\omega^{1}\omega^{4}(1\pm\omega^{2}\omega^{3})\cdot\frac{T_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}+o(1),$

### v3

@@ 1 Introduction
While most analyses of the CHSH inequality rely on probabilistic or measure-theoretic formulations, comparatively little attention has been paid to how the availability and timing of information themselves constrain possible correlations. The objective of this paper is to shed new light on the conditions under which the CHSH inequality holds or is violated, from this informational perspective. To this end, this paper employs the framework of game-theoretic probability theory [13, 14].
@@ 4.1 A hidden-variable game with the locality loophole
  Hidden-variable game with the locality loophole   Players : Scientist A, Scientist B, Nature A, and Nature B. Protocol : $F_{0}^{A}=F_{0}^{B}=1$. FOR $n\in\mathbb{Z}_{>0}$ : Scientist A announces $\{q_{n}^{A}(\,\cdot\,\mid s)\in\mathcal{P}(\Omega)\}_{s\in\Theta^{A}}$ and $s_{n}\in\Theta^{A}$. Nature A announces $\omega_{n}^{s_{n}}\in\Omega$. Scientist B announces $\{q_{n}^{B}(\,\cdot\,\mid\omega_{n}^{s_{n}},s_{n},t)\in\mathcal{P}(\Omega)\}_{t\in\Theta^{B}}$ and $t_{n}\in\Theta^{B}$. Nature B chooses $\lambda_{n}\in X_{s_{n}}^{-1}(\{\omega_{n}^{s_{n}}\})$ and announces $\omega_{n}^{t_{n}}:=X_{t_{n}}(\lambda_{n})$. $F_{n}^{A}:=F_{n-1}^{A}\cdot\frac{q_{n}^{A}(\omega_{n}^{s_{n}}\mid s_{n})}{p(\omega_{n}^{s_{n}}\mid s_{n})}$, $F_{n}^{B}:=F_{n-1}^{B}\cdot\frac{q_{n}^{B}(\omega_{n}^{t_{n}}\mid\omega_{n}^{s_{n}},s_{n},t_{n})}{p(\omega_{n}^{t_{n}}\mid\omega_{n}^{s_{n}},s_{n},t_{n})}$. END FOR.
@@
In the hidden-variable game with the locality loophole, Scientists A and B have a joint strategy that ensures $\lim_{n\to\infty}F_{n}^{A}=\infty$ or $\lim_{n\to\infty}F_{n}^{B}=\infty$ unless, for any $s\in\Theta^{A}$, $t\in\Theta^{B}$ and $a,b\in\Omega$,
@@ 4.2 A hidden-variable game with loopholes closed
$U_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4};s,t):=$ $\left|\left\{k\middle|\begin{aligned} \;&1\leq k\leq n,\,s_{k}=s,\,t_{k}=t,\\ &(X_{1}(\lambda_{k}),X_{2}(\lambda_{k}),X_{3}(\lambda_{k}),X_{4}(\lambda_{k}))=(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\end{aligned}\right\}\right|,$
$V_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4}):=$ $\left|\left\{k\middle|\begin{aligned} \;&1\leq k\leq n,\,\\ &(X_{1}(\lambda_{k}),X_{2}(\lambda_{k}),X_{3}(\lambda_{k}),X_{4}(\lambda_{k}))=(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\end{aligned}\right\}\right|$
$=$ $\sum_{(s,t)\in\Theta}U_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4};s,t),$
$W_{n}(s,t):=$ $|\{k\mid 1\leq k\leq n,\,s_{k}=s,\,t_{k}=t\}|$
$=$ $\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}U_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4};s,t).$
The quantity $U_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4};s,t)$ counts how often the hidden variable assignments produce the outcome quadruple $(\omega^{1},\omega^{2},\omega^{3},\omega^{4})$ under the measurement setting $(s,t)$. The quantities $V_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})$ and $W_{n}(s,t)$ are the corresponding marginal counts. We then define empirical distributions on $\Omega^{4}\times\Theta$, $\Omega^{4}$, and $\Theta$ by
$\hat{P}_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4};s,t)$ $:=\frac{U_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4};s,t)}{n},$
$\hat{Q}_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})$ $:=\frac{V_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n},$
$\hat{R}_{n}(s,t)$ $:=\frac{W_{n}(s,t)}{n}.$
@@
  Hidden-variable game with loopholes closed   Players : Scientist AB, Nature AB. Protocol : $F_{0}^{AB}=1$, $I_{0}^{AB}=1$. FOR $n\in\mathbb{Z}_{>0}$ : Scientist AB announces $\{(q_{n}^{AB}(\,\cdot,\cdot\,\mid s,t))\in\mathcal{P}(\Omega^{2})\}_{(s,t)\in\Theta}$ and $(s_{n},t_{n})\in\Theta$. Nature AB announces $\lambda_{n}\in\Lambda$ and $\omega_{n}^{s}:=X_{s}(\lambda_{n})$ for $s\in\{1,2,3,4\}$. $F_{n}^{AB}:=F_{n-1}^{AB}\cdot\frac{q_{n}^{AB}(\omega_{n}^{s_{n}},\omega_{n}^{t_{n}}\mid s_{n},t_{n})}{p(\omega_{n}^{s_{n}},\omega_{n}^{t_{n}}\mid s_{n},t_{n})}$. $I_{n}^{AB}:=I_{n-1}^{AB}\cdot\frac{\hat{P}_{n}(\omega_{n}^{1},\omega_{n}^{2},\omega_{n}^{3},\omega_{n}^{4};s_{n},t_{n})}{\hat{Q}_{n}(\omega_{n}^{1},\omega_{n}^{2},\omega_{n}^{3},\omega_{n}^{4})\cdot\hat{R}_{n}(s_{n},t_{n})}$. END FOR.
The second capital process $I_{n}^{AB}$ is designed to test, in a sequential manner, whether the empirical joint distribution on $\Omega^{4}\times\Theta$ factorizes asymptotically into the product of its marginals. It is defined recursively via a likelihood-ratio-type update between $\hat{P}_{n}$ and $\hat{Q}_{n}\cdot\hat{R}_{n}$, so that any systematic dependence between Nature’s hidden-variable declarations and the chosen measurement settings leads to the divergence of $I_{n}^{AB}$.
Structurally, the present game differs from the locality-loophole game in only two respects. First, Scientist AB introduces the additional capital process $I_{n}^{AB}$ in order to operationally enforce a measurement-independence-like constraint. Second, by changing the timing of the announcements (relative to the locality-loophole game), we reformulate the dependence between the hidden variable $\lambda_{n}$ and the measurement settings $(s_{n},t_{n})$ from a purely temporal dependence into a dependence that is empirically testable through frequencies. Although the timing of moves differs from that in the locality-loophole game, this does not restrict Nature’s ability to generate outcome pairs $(\omega_{n}^{s_{n}},\omega_{n}^{t_{n}})$ as functions of $(s_{n},t_{n})$. The essential difference is that the loophole-closed game makes such dependence detectable within the game, which ultimately leads to the following incompatibility result.
@@
In the hidden-variable game with loopholes closed, Scientist AB has a strategy $q:\Omega^{\ast}\rightarrow\mathcal{P}(\Omega^{2})$, that ensures $\lim_{n\to\infty}F_{n}^{AB}=\infty$ unless, for any $(s,t)\in\Theta$ and $a,b\in\Omega$,
@@
In addition, Scientist AB’s second capital process $I_{n}^{AB}$ diverges, $\lim_{n\to\infty}I_{n}^{AB}=\infty$, unless, for any $(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}$ and $(s,t)\in\Theta$,
@@
Furthermore, Nature AB cannot satisfy (1) and (2) simultaneously. Thus in this game, either $\lim_{n\to\infty}F_{n}^{AB}=\infty$ or $\lim_{n\to\infty}I_{n}^{AB}=\infty$ must occur.
@@
Theorem 4.2 states, in game-theoretic terms, that Scientist AB possesses a winning strategy in the hidden-variable game with loopholes closed. More precisely, Scientist AB can choose a strategy $q:\Omega^{\ast}\to\mathcal{P}(\Omega^{2})$ such that Nature AB cannot keep both capital processes $F_{n}^{AB}$ and $I_{n}^{AB}$ bounded. Scientists run two capital processes in parallel.
  • $F_{n}^{AB}$ tests whether the empirical conditional frequencies converge to the CHSH table.
  • $I_{n}^{AB}$ tests whether the empirical distribution on $\Omega^{4}\times\Theta$ factorizes asymptotically into the product of its marginal distributions on $\Omega^{4}$ and on $\Theta$, providing a frequency-based analogue of measurement independence.
Note that (2) is weaker than the probabilistic independence assumptions usually imposed in the context of the CHSH inequality. Even under this weaker requirement, Nature must either reproduce the CHSH table at the level of empirical conditional frequencies (1) or exhibit setting-independence-like behavior, in the frequency sense (2), but it cannot achieve both simultaneously. Thus, as $n\to\infty$, at least one of $F_{n}^{AB}$ or $I_{n}^{AB}$ necessarily diverges.
At first sight this may seem at odds with the empirical situation: in actual CHSH experiments, Bell-type inequalities are violated in accordance with quantum-mechanical predictions. The point, however, is that the theorem is not merely saying that “a violation occurs.” Rather, it clarifies what must give way within a local hidden-variable (LHV) picture once the relevant loopholes are controlled. Physically, an LHV model can survive asymptotically only by abandoning one of the two: either it fails to reproduce the CHSH table, or it gives up measurement independence by correlating the setting choices with the hidden variable.
Operationally, the two capital processes function as sequential test statistics that diagnose compatibility with each requirement along the data stream. Thus, in a loophole-closed setting, Scientists obtain an operational winning strategy: they inevitably detect a breakdown, either in reproducing the CHSH table at the level of empirical frequencies or in setting-independence-like behaviour, thereby excluding LHV explanations in a data-driven sense.
@@ A Proofs of Theorems presented in Section 4
We divide the proof into three parts. First, as a direct consequence of Theorem 2.1, Scientist AB has a strategy $q:\Omega^{\ast}\rightarrow\mathcal{P}(\Omega)$, that ensures $\lim_{n\to\infty}F_{n}^{AB}=\infty$ unless, for all $(s,t)\in\Theta$ and $a,b\in\Omega$,
@@
We note that as in Theorem 4.1, this step is merely an application of Theorem 2.1, which forces Nature to conform to the CHSH table at the level of empirical frequencies. The novelty of Theorem 4.2 lies not in this enforcement itself, but in the fact that its consequence is incompatible with the boundedness of the capital process $I_{n}^{AB}$.
Second, we evaluate the capital process $I_{n}^{AB}$. To simplify the symbols, we write $\tau=(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}$ and $u=(s,t)\in\Theta$. Then by the recursion formula of $I_{n}^{AB}$,
$I_{n}^{AB}=\frac{n!\cdot\prod_{\tau\in\Omega^{4}}\prod_{u\in\Theta}U_{n}(\tau;u)!}{\prod_{\tau\in\Omega^{4}}V_{n}(\tau)!\cdot\prod_{u\in\Theta}W_{n}(u)!}.$
@@
$\frac{\log I_{n}^{AB}}{n}=$ $\frac{1}{n}\left\{\log n!+\sum_{\tau\in\Omega^{4}}\sum_{u\in\Theta}\log U_{n}(\tau;u)!-\sum_{\tau\in\Omega^{4}}\log V_{n}(\tau)!-\sum_{u\in\Theta}\log W_{n}(u)!\right\}$
$=$ $\log n+\sum_{\tau,u}\frac{U_{n}(\tau;u)}{n}\log U_{n}(\tau;u)$
$-\sum_{\tau}\frac{V_{n}(\tau)}{n}\log V_{n}(\tau)-\sum_{u}\frac{W_{n}(u)}{n}\log W_{n}(u)+O\left(\frac{\log n}{n}\right)$
$=$ $\sum_{\tau,u}\frac{U_{n}(\tau;u)}{n}\left(\log\frac{U_{n}(\tau;u)}{n}-\log\frac{V_{n}(\tau)}{n}-\log\frac{W_{n}(u)}{n}\right)+O\left(\frac{\log n}{n}\right)$
@@
then $\limsup I_{n}^{AB}=\infty$.
@@
$\frac{U_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4};s,t)}{W_{n}(s,t)}-\frac{V_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}$
@@
$\frac{\widetilde{U_{n}}(\omega^{s},\omega^{t};s,t)}{W_{n}(s,t)}-\frac{\widetilde{V_{n}}(\omega^{s},\omega^{t})}{n}=o(1),$ (A.2)
@@
$\widetilde{U_{n}}(\omega^{s},\omega^{t};s,t)$ $:=|\{k\mid 1\leq k\leq n,\,X_{s}(\lambda_{k})=\omega^{s},\,X_{t}(\lambda_{k})=\omega^{t},\,s_{k}=s,\,t_{k}=t\}|,$
$\widetilde{V_{n}}(\omega^{s},\omega^{t})$ $:=|\{k\mid 1\leq k\leq n,\,X_{s}(\lambda_{k})=\omega^{s},\,X_{t}(\lambda_{k})=\omega^{t}\}|.$
@@
$C_{n}(s,t):=\sum_{(\omega^{s},\omega^{t})\in\Omega^{2}}\omega^{s}\omega^{t}\cdot\frac{\widetilde{U_{n}}(\omega^{s},\omega^{t};s,t)}{W_{n}(s,t)}.$
@@
$=\sum_{(\omega^{1},\omega^{3})\in\Omega^{2}}\omega^{1}\omega^{3}\cdot\frac{\widetilde{U_{n}}(\omega^{1},\omega^{3};1,3)}{W_{n}(1,3)}-\sum_{(\omega^{1},\omega^{4})\in\Omega^{2}}\omega^{1}\omega^{4}\cdot\frac{\widetilde{U_{n}}(\omega^{1},\omega^{4};1,4)}{W_{n}(1,4)}$
$=\sum_{(\omega^{1},\omega^{3})\in\Omega^{2}}\omega^{1}\omega^{3}\cdot\frac{\widetilde{V_{n}}(\omega^{1},\omega^{3})}{n}-\sum_{(\omega^{1},\omega^{4})\in\Omega^{2}}\omega^{1}\omega^{4}\cdot\frac{\widetilde{V_{n}}(\omega^{1},\omega^{4})}{n}+o(1)$
$=\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}\omega^{1}\omega^{3}\cdot\frac{V_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}-\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}\omega^{1}\omega^{4}\cdot\frac{V_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}+o(1)$
$=\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}\omega^{1}\omega^{3}(1\pm\omega^{2}\omega^{4})\cdot\frac{V_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}$
$\ \ -\sum_{(\omega^{1},\omega^{2},\omega^{3},\omega^{4})\in\Omega^{4}}\omega^{1}\omega^{4}(1\pm\omega^{2}\omega^{3})\cdot\frac{V_{n}(\omega^{1},\omega^{2},\omega^{3},\omega^{4})}{n}+o(1),$